OpenStudy (anonymous):

A security light is being installed outside a loading dock. The light must be placed at a 52° angle so that it illuminates a parking lot. If the distance from the end of the parking lot to the loading dock is 100 feet, what is the height of the security light? A. 50 feet B. 39.06 feet C. 100 feet D. 78.13feet

OpenStudy (anonymous):

Security Light

OpenStudy (anonymous):

If we let the height of the light be "x", then: tan 52 = 100 / x -> x = 100 / (tan 52)

OpenStudy (anonymous):

78

OpenStudy (anonymous):

tan 52 = 1.279941632193078780311029847572 so, x = 100 / (tan 52) -> x = 100 / 1.279941632193078780311029847572 x = 78.128562650671739706294997196227 Good job again! @Joshuam1

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

new post